OpenStudy (anonymous):

Determine the value of x in the equation 12x + 4= 46

Directrix (directrix):

12x + 4= 46 12x = 42 x = 42/ 12 = 21/6

Directrix (directrix):

x = 42/ 12 = 21/6 = 7/2

OpenStudy (anonymous):

12x+4=46 12x=46-4 12x=42 x=42/12 x=3.5 you can check that by putting 3.5 in place of x (12*3.5+4=46)

OpenStudy (anonymous):

thx!!!!!!!